# Massimiliano Ciranni1 These authors contributed equally to this work.    Vito Paolo Pastore1,2∗    Roberto Di Via1∗    Enzo Tartaglione3    Francesca Odone1    Vittorio Murino2,4

### v2

@@ Abstract
Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data which, whenever containing strong spurious correlations between specific attributes and target labels, can result in unrecoverable biases in model predictions. Tackling these biases is crucial in improving model generalization and trust, especially in real-world scenarios. This paper presents Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods in model debiasing while exploiting the inherent bias-learning tendency of diffusion models. Our approach leverages conditional diffusion models to generate synthetic bias-aligned images, used to train a bias amplifier model, to be further employed as an auxiliary method in different unsupervised debiasing approaches. Our proposed method, which also tackles the common issue of training set memorization typical of this type of techniques, beats current state-of-the-art in multiple benchmark datasets by significant margins, demonstrating its potential as a versatile and effective tool for tackling dataset bias in deep learning applications.
@@
Figure 1: Diffusing DeBias (DDB) Snapshot. A conditional diffusion model learns per-class distributions from a biased dataset, resulting in synthetic generations that reflect these biases (referred to as the bug). These are used to train a Bias Amplifier, converting the bug into a feature useful for debiasing recipes to mitigate bias.
@@ 1 Introduction
In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives. Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]). As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts. To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5]. Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6]. However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5]. Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10]. Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
@@
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15]. However, we claim that what appears to be a bug for generative tasks can be turned into a potential feature for image classification model debiasing (see Fig. 1), as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes. For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness, we incorporate DDB into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches outperform the state-of-the-art on popular synthetic and real-world benchmark datasets.
To summarize, our main contributions can be summarized as follows:
  • We introduce DDB, a novel unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information. The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec. 3).
@@ 2 Related Work
Various bias mitigation approaches have been explored in the last few years. In the following, we will address the two general setups, supervised and unsupervised, and we finally illustrate how bias is considered in the context of generative diffusion models.
@@
Supervised Debiasing. It refers to methods exploiting available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17]. Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8]. At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Typical approaches can be distinguished between two-step methods (where the first step identifies bias in training data, generally through pseudo-labeling [11, 9, 4]), and end-to-end (where bias mitigation is performed online with dedicated optimization objectives [6, 20]). An important limitation of current unsupervised methods is avoiding training bias-conflicting overfitting without a properly balanced validation set (a-priori not available), still a major argument of discussion among the community. Employing techniques such as ensembling and noise-robust losses have been proposed, though requiring careful tuning of all the hyperparameters which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. [6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21]. At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. Jung et al. [22] introduced a debiasing method where a GAN [23] is trained for style-transfer between several biased appearances. A target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations. In [11], a model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In [1], a set of bias-amplified classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to detect and mitigate biases in classification models [24, 25, 26].
Diffusion Models and Bias. Diffusion models have emerged as powerful generative models capable of producing high-quality data [27]. However, like many deep learning models, they are exposed to biases present in the training data or introduced during the modeling process [15], which can manifest as distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [28]. Several recent studies tackle bias in diffusion models through different approaches, e.g., by exposing and correcting hidden biases in the diffusion process [29], introducing dual-training frameworks for more robust and fair generation [14], addressing specific biases in facial image generation [15], or proposing training methods that counter biased input data effects [13]. While these studies focus on mitigating the bias in diffusion model training aiming to produce fairer and more accurate results, our proposed solution frames this inherent behavior in a novel perspective.
Unlike previously cited works, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class. This side-effect allows us to sample new synthetic images, sharing the same bias patterns, and use them for unsupervised model debiasing, hence avoiding many limitations of the current state of the art by construction. Actually, exclusively using synthetic data in auxiliary model training prevents overfitting on original real data, while also resulting in a fully-fledged bias amplified model to be exploited for bias identification.
3 Approach
@@
This Section illustrates the proposed debiasing approach, schematically depicted in Figure 2. First, we provide a general description of the problem setting (Sec.  3.1). Second, we report a detailed description of the DDB’s two main components, which include bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.2).
@@ 3.1 Problem Setting
Let us consider a dataset of images and labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$, and assume the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors (e.g., context, appearance, acquisition noise). In this case, it will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases. In these conditions, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or not annotated in a realistic setting [13]. If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization, due to the strong correlation between $b$ and $y$ (often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 30]), dominating over the true target distribution semantics.
Data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12]. For instance, given a CDPM modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we can expect that its generations will be biased [12, 13]. While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e.$\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.

### v3

# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

@@ Abstract
Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper aims at facing this problem by leveraging bias amplification with generated synthetic data: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which replace the original training set for learning an effective bias amplifier model that we subsequently incorporate into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples memorization in learning auxiliary models, typical of this type of techniques, our proposed method beats current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
@@
Figure 1: Comparison of bias amplifier training approaches. (a) Traditional methods that use real biased datasets lead to the memorization of biased-conflicting samples, resulting in suboptimal auxiliary models. (b) Our proposed approach leverages conditional diffusion models to learn class-specific biases and amplify them into synthetic images. Such generations can replace the original training set for learning an effective Bias Amplifier, eliminating memorization effects by avoiding exposure to real data.
@@ 1 Introduction
In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives. Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]). As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5]. Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6]. However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5]. Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10]. Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
@@
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully biased synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15]. However, we claim that what is considered a problem for generative tasks can be turned into a feature for image classification model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization, by construction (see Figure 1). For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches outperform the state-of-the-art on popular benchmark datasets. To summarize, our main contributions can be summarized as follows:
  • We introduce DDB, a novel and effective unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information. The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec. 3).
@@ 2 Related Work
In this section, we provide a broad categorization of model debasing works.
@@
Supervised Debiasing. Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17]. Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8]. At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process. We categorize these methods into two-step  [11, 9, 4]), and end-to-end  [6, 20]), in the following paragraphs.
End-to-end approaches. Here, bias mitigation is performed online with dedicated optimization objectives. Employing techniques such as ensembling and noise-robust losses have been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. [6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21]. At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. In DebiAN, Li et al. [20] jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
Two-step methods. Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists in actual bias mitigation exploiting the inferred pseudo-labels. In [11], the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In [1], a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models [22, 23, 24].
Among the described methods, a subset exploit the auxiliary model for bias amplification ([5, 6, 9]) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples, with final debiasing performance strongly dependent on such behavior, and specifically, on how strongly the auxiliary model captures training bias [3, 5]. However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs [10, 3], especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in [8, 25]. One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, this being en emerging argument of discussion among the community [10, 3].
For this reason, in this work and differently from the previously cited research, we propose that syntethic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction. However, an unsupervised real-world scenario poses a problem for generating a pure distribution of bias-aligned samples. As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier, inspired by recent works, showing how they can manifest distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [26], when exposed to biases present in the training data  [15]. While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation [27, 28], up to our knowledge, we are the first to propose the usage of generative models for bias amplification. The closest available work, employing generative models for model debiasing, is Jung et al. [29], where the authors introduce a debiasing method exploiting a GAN [30] trained for style-transfer between several biased appearances, with a different objective. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
In our work, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of original train data. Our method provides a memorization-free bias amplification model, capable of providing precise supervisory signals for accurate model debiasing virtually compatible with any debiasing approach involving auxiliary models.
Figure 2: Schematic representation of our DDB framework. The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract supervisory bias signals from real images. These signals are used to guide the training process of a target debiased model by designing two debiasing recipes (i.e., 2-step and end-to-end methods).
3 The Approach
@@
Our proposed debiasing approach is schematically depicted in Figure 2. In this section, we provide at first a general description of the problem setting (Sec.  3.1), and then, we illustrate in detail DDB’s two main components, which include bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.3).
@@ 3.1 Problem Setting
Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution. Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e.possible biases such as context, appearance, acquisition noise, viewpoint, etc.. In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases. In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting [13] In this context, biased and unbiased samples equivalently refer to bias-aligned and bias conflicting samples.. If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization. This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 31]), which is dominating over the true target distribution semantics.
It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12]. For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in [12, 13]. While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.